Find the value of b that makes the system of equations have the solution (3,5)
y=3x-4
y=bx+2

Found 2 solutions by dabanfield, richwmiller:
Answer by dabanfield(803)

y=bx+2
If (3,5) is a solution then we must have:
5 = b*3 + 2
5 - 2 = 3*b
3 = 3*b
b = 3/3
b = 1

Answer by richwmiller(17219)

Since 3,5 will be a solution for both equations, we can place it in either or both.
y=bx+2
5=3b+2
3=3b
1=b answer
check
5=3*3-4
5=9-4
ok
5=1*3+2
5=5
ok
Answer b=1